OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 6m = 4 + 4n Step 1: −6(m) = −6(8 + 2n) [Equation 1 is multiplied by −6.] 6m = 4 + 4n [Equation 2] Step 2: −6m = −48 − 12n [Equation 1 in Step 1 is simplified.] 6m = 4 + 4n [Equation 2] Step 3: −6m + 6m = −48 − 12n + 4n [Equations in Step 2 are added.] Step 4: 0 = −48 − 8n Step 5: n = −6 In which step did the student first make an error? Step 3 Step 5 Step 4 Step 2 I chose D? Am I correct ?

OpenStudy (anonymous):

Yes I believe so

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

No problem any more questions just tag me

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

NO it was step 3

OpenStudy (anonymous):

oh ok

OpenStudy (triciaal):

add again

OpenStudy (triciaal):

by now you should notice that not everyone has the correct answer so you need to understand what is expected to solve the problem

OpenStudy (anonymous):

okay

OpenStudy (texaschic101):

let me check this...it looks confusing to me multiply eq 1 by -6 -6m = -48 - 12n add the 2 equations -6m = -48 - 12n 6m = 4 + 4n --------------add 0 = - 44 - 8n <====step 4 is wrong

OpenStudy (triciaal):